OpenStudy (anonymous):

Solve by Substitution Method. 2x-5y=-1 x - 3y=-3

OpenStudy (anonymous):

From equation 2, x= -3 + 3y Substitute this into equation 1, we get 2(-3+3y)-5y=-1 -6+6y-5y=-1 y=-1+6 y=5

OpenStudy (anonymous):

Substituting y=5 back into equation 2, we get x - 3(5) = -3 x - 15 = -3 x = -3 + 15 x =12

OpenStudy (anonymous):

Therefore solutions are, x=12 , y=5

OpenStudy (anonymous):

Thanks a lot